Assume a 170-W loudspeaker broadcasts sound equally in all directions and produces sound with a level of 108 dB at a distance of 1.28 m from its center.

(a) Find its sound power output. If a salesperson claims the speaker is rated at 170 W, he is referring to the maximum electrical power input to the speaker. W

(b) Find the efficiency of the speaker, that is, the fraction of input power that is converted into useful output power. %
Answer #1

A)

Sound intensity, B = 10 log(I/Io)

108 = 10 log(I/10^-12)

I = 0.063 W/m^2

Power, P = I x A = 0.063 x 4 pi x (1.28^2) = 1.30 W

B)

Efficiency, n = P/P' = 1.3/170 = 0.76%

Comment in case any doubt please rate my answer ..